17 July 2017
Monday 17th of July 2017

Helle, Khanghyun and Mayu making snow core measurements using the portable CFA system.


Today was a day of filtering in the drill trench. The drillers estimate that some 180 kg of slush had been retrieved from the bore hole in total by the end of the day, and the bottom of the bore hole has still not been reached by the filter. Hence filtering will continue tomorrow.

On the surface the normal scientific program continued, but there was also time for Helle, Mayu and Khanghyun to do experimental measurements with the portable CFA box, a total of 5 m of snow cores went through the system and annual cycles could be seen in the conductivity measurements.

In the science trench physical properties measurements continued as planned. Furthermore documenting, packing and strapping of ice core boxes were commenced in anticipation of this week’s flight. The AWI drill system was also dismounted and made ready for packing.

What we did today:

  1. Continued filtering of bore hole, some 180 kg of slush retrieved so far.
  2. Started pack down of AWI drill system.
  3. Started documentation and strapping of ice core boxes.
  4. Measured a total of 5 meters of snow cores with portable CFA system.
  5. Water vapour, methane, aerosols, snow and boundary layer measurements
  6. Physical properties measurements.
  7. Water isotope sampling for sublimation studies.
  8. Mounting end stops on core throughs.
  9. Removing snow drifts with snow blower at trench entrances and tidying up in camp after snow.

Weather: Mostly cloudy with snow showers in the evening. Wind from SW turning to NE during the day, up to 6 m/s. Temperature between -20°C and -8°C.

FL, Bo Vinther

← Previous entry   Next entry →